find an equation for the line parallel to the 4x+6y=-7 having the same y-intercept as 2x-5y=-5
Answer by mananth(16946)

2x-5y=-5
5y=2x+5
y=(2/5)x+1
the y intercept = +1
4x+6y=-7
6y=-4x-7
y=(-4/6)x-7/6
y=(-2/3)x=7/6
The slope of the required line is (-2/3) & y intercept = 1
So the equation will be y=(-2/3)x+1
m.ananth@hotmail.ca